Question 4: Debt markets (15 Marks)

Please answer the following questions. Show all your workings when calculations are required and round off your FINAL result to TWO decimal places

a) Pearson Publishing Ltd needs to borrow money for two purposes: purchase of inventory and purchase of a building to expand its business. Please advise this company on how to raise funds for these two purposes. In your discussion you need to define and distinguish between the debt markets advised. (6 marks)

b) Pearson decides to issue a 90-day commercial paper at a yield of 2.5% p.a. If the face value of the paper is $150,000, how much fund will Pearson be able to raise? (3 marks)

c) The company has decided to issue a bond at a Face Value of $1,000. The coupon rate of the bond is 8% per annum and the coupons are paid semi-annually. The bond has a term to maturity of 5 years and the yield-to-maturity is 8% per annum as well. Without calculation, briefly explain what the implied price of this bond would be? (2 marks)

d) If an investor bought the bond at issuance at the implied price in part c) and then sold it at $1030 after holding it for 1.5 years, what was his holding period yield per annum? (4 marks)
Answer #1

Ans-(a)

Funding Inventory is a short term requirement and this can be funded through Commercial Paper.These are unsecured short term instruments issued at discount in face value and reflect the prevailing market interets rate.Pearson can take use of this instrument but ofcourse subjected that the credit ratings are good.Alternatively it can also take short term loan from Bank.

Funding or expansion is long term scenario and hence.Pearson can issue bonds or debentures.Thes carry a coupon rate payable periodically to the lenders and has a longer maturity period.

Ans-(b)

The price of the bond to get a yield of 2.5% can be calculated using the following formula

Price = Face Value/ [1 + yield x (no. of days to maturity/365)]

FV = 150000

Yield = 2.5%

No. of days to maturity = 90

Piutting the above in the formula :

Price = $ 149081

Ans-(c)

Implied price of the bond will be the sum of the Present value of all the interest payments and the maturity payment.The Present value to be calculated using the YTM rate of 8% and not the coupon rate.

YTM is the total return on a bond if held till maturity.It is considered the IRR of a bond

Ans-(d)

The PV of the Bond will be same as Face value since Coupon rate and market rate are both 8%

Total Return for the Investor:

3 Interest payments of 40 each i.e. 1000*8%/2 = 40 (since semiannual) = 40*3=$120

Gain in Market Price = $1030 - $1000 = $30

Total Gain = 120 +30 = 150

Holding Period Rate (HPR) = 150/1000*100 = 15% 0r (1115-1000)/1000 = 115%

Annualized HPR = ((1+1.15)^1/3) -1 = 28%